• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 09:55:50 on March 06, 2012

    Entry number 364228

    This entry is a followup to: 363823

    keyword=harp test plots

    Alexandre posted the harp test plots on Elog 1682635. See also later plot Elog 168876. (In the future, please have the critical information also posted on Alog or post a link to the Elog entry.) Signals at 1H05A improved with increased current. At the highest current tested, 8 uA (pulsed), they are visible, but still marginal.